# Noise2Ghost: Self-supervised deep convolutional reconstruction for ghost imaging

### v2

@@ Abstract
We present a new self-supervised deep-learning-based Ghost Imaging (GI) reconstruction method, which provides unparalleled reconstruction performance for noisy acquisitions among unsupervised methods. We present the supporting mathematical framework and results from theoretical and real data use cases. Self-supervision removes the need for clean reference data while offering strong noise reduction. This provides the necessary tools for addressing signal-to-noise ratio concerns for GI acquisitions in emerging and cutting-edge low-light GI scenarios. Notable examples include micro- and nano-scale x-ray emission imaging, e.g., x-ray fluorescence imaging of dose-sensitive samples. Their applications include in-vivo and in-operando case studies for biological samples and batteries.
journal: opticajournal
@@ 1 Introduction
Thanks to all these advantages, GI has the potential to spark many breakthroughs in applications in various fields ranging from biomedical imaging to remote sensing. However, GI has only seen limited applicability so far, especially in the above-mentioned nano-scale imaging applications. In these cases, the observed signals are affected by overwhelmingly high Poisson noise levels due to their very small photon fluxes. This noise type does not enjoy the intrinsic noise reduction of GI, known as Felgett’s advantage, thus reducing GI’s advantage over PB [8]. Here, we propose a GI reconstruction method particularly suited for working with acquisitions affected by high noise levels, regardless of whether the noise type is Gaussian or Poisson. Our method leverages the latest developments in unsupervised machine learning (ML) techniques, so it does not require high-quality reference data. This work is a step towards practical applications of GI, particularly for photon-limited scenarios like nano-scale x-ray emission imaging [10].
@@ 2.1 Existing methods
GI’s ability to obtain images with more pixels than the acquired realizations means that the associated reconstruction problem is undersampled. These undersampled reconstructions require some prior knowledge (regularization) to recover an accurate image: Unregularized least-squares (LS) reconstructions present specific high-frequency large-scale artifacts that resemble high noise levels. Methods employing convolutional neural networks (CNNs) are currently the best-performing tools for image processing applications. They can capture small- and large-scale correlations and learn the most suitable representations to capture the features present in the treated signals (acting like regularizers) [12]. Thus, they are also better suited for GI reconstructions than traditional variational methods, e.g., Total Variation (TV) minimization, which use strong but also rather simple a priori assumptions on the types of features present in the reconstructed images.
@@
ML methods and CNN architectures have both seen an explosion in development and applications over the last decade. Quite a few ML methods have been specifically developed for GI reconstructions. Notable examples are supervised methods [13], where a CNN (e.g., a U-net [14]) compares examples of LS GI reconstructions from highly undersampled acquisitions against their corresponding high-quality reconstructions. The network learns to identify the corrupted features in the LS images and to replace them with their expected shape. This approach is heavily dependent on large amounts of known examples that present strong similarity to the images of interest. It also requires specific training data for each acquisition parameter variation (e.g., noise levels, image size, levels of undersampling, etc).
@@
Untrained generator-network methods do not require pre-training of the model against reference high-quality images, removing the need for large databases of reference images [15, 16]. They can incorporate the knowledge of the image formation model in the learning algorithm (an approach of physics-informed ML), and make the trained model learn to produce an image that fits the acquired data. A good example is the method called GIDC [17], which is based on the deep image prior (DIP) [12]. As a drawback, these methods exhibit lower reconstruction quality for high noise levels in the acquired data. They delegate noise reduction solely to regularization and early stopping. Similar to DIP-based methods, implicit neural representations (INRs) -based methods use untrained NNs to reconstruct an image without high-quality reference data [18, 10]. INR-based methods use multi-layer perceptron (MLP) NNs and accept pixel coordinates instead of input images. During the training, the MLP learns a representation of the imaged object, which produces the desired image for the input coordinates.
@@
Self-supervised methods like Noise2Noise (N2N) use the same training procedure as the supervised methods, but they use noisy images as targets [19]. They only require two noisy realizations of the same measurement (e.g., two noisy images of the same scene) and exploit the fact that the only difference between the two measurements should be the noise. Thus, given one of the two measurements, they train the CNN to predict the other and vice versa. As previously mentioned, CNNs can capture small and large-scale correlations, leading them to learn the features of the measurements while discarding their uncorrelated parts, i.e., the noise. N2N can be adapted to work with inverse problems like computed tomography [20]. However, as we will also see in Subsec.˜2.5, it does not handle the so-called missing realization problem, which refers to artifacts arising from not having enough realizations to represent the imaged objects, and that is usually an important source of noise in GI reconstructions.
@@ 2.2 Forward model and variational methods
$\begin{split}&\mathbb{E}_{y}\lVert WN_{\theta}(r)-y\rVert_{2}^{2}\\ &=\mathbb{E}_{b,\epsilon}\lVert WN_{\theta}(W^{\dagger}(b+\epsilon))-(b+\epsilon)\rVert_{2}^{2}\end{split}$ (4)
@@ 2.3 Proposed method: Noise2Ghost (N2G)
We partition the realizations in $k$ splits $k\in[1,K]$, producing $K$ different sub-reconstructions (i.e., a k-tuple of reconstructions). Each sub-reconstruction in the said k-tuple has unique noise characteristics compared to the other reconstructions, and it provides different information on the FoV. From the partitioned data vectors $y_{k}$ we compute the sub-reconstructions $x_{k}$ as follows (Fig.˜2a):
@@
Each sub-reconstruction $x_{k}$ represents the same signal but has a different noise. In N2G, we feed each of these sub-reconstructions to the model, and then we optimize the following minimization problem (Fig.˜2b) by comparing them against the realizations that were not used to reconstruct them:
@@
Since the different sub-reconstructions $x_{k}$ should represent the same signal, we would like the model to produce the same result for each. Thus, the model should learn to differentiate between the actual signal and random noise. This is in opposition to Eq.˜3a, where the same realizations were used for the input reconstruction and as targets in the learning loss. Finally, Fig.˜2c represents how Eq.˜6b produces the reconstructed images. In Sec. S4 of the Supplementary Information We briefly discuss the machine learning models, in particular neural networks (NN), that N2G could leverage.
@@ 2.4 Self-supervised random noise reduction
$\begin{split}\mathbb{E}_{b,\epsilon,i\neq k}&\lVert W_{i}N_{\theta}(x+v_{k}+W^{\dagger}\epsilon_{k})-b_{i}\rVert_{2}^{2}\\ &+\mathbb{E}_{b,\epsilon,i\neq k}\lVert b_{i}-(b_{i}-\epsilon_{i})\rVert_{2}^{2}.\end{split}$ (9)
This can be further simplified into:
$\mathbb{E}_{b,i\neq k}\lVert W_{i}N_{\theta}(x_{k})-b_{i}\rVert_{2}^{2}+\mathbb{E}_{\epsilon,i\neq k}\lVert\epsilon_{i}\rVert_{2}^{2}$ (10)
@@ 2.5 Relationship with Noise2Inverse
where $x_{i}=W_{i}^{\dagger}y_{i}$ is the pseudo-inverse (LS reconstruction) of the partial set of masks $W_{i}$ and corresponding buckets $y_{i}$ with $i\in[1,K]$. Theoretical grounds support this formulation, and it should retrieve the same result as Eq.˜6, regarding its noise reduction aspect. However, it does not address the missing realization artifacts and suffers from strong technical limitations. In ghost imaging, artifacts arising from both missing realization noise and random noise are structured and have long-range correlations, i.e., they generally cover the whole FoV. Commonly used models like CNNs have a limited receptive field (i.e., the region around each pixel that the CNN probes and that provides the context for the said pixel) that may not extend to the entire FoV. Depending on the model, this could have a strong negative impact on the denoising performance.
More precisely, about the missing realization artifacts, in [22] it is stated that Eq.˜11 is unable to cope with them. The LS sub-reconstructions cope with the missing realizations by acting as if their bucket values were set to 0. This selects one specific solution $v_{k}$ for each $x_{k}$, out of the infinite solutions in $\ker W_{k}$ that satisfy Eq.˜1. This specific choice of $v_{k}$ is probably incorrect, but is imposed by N2I as the learning target in Eq.˜11. The model learns to discard the inconsistent features of different $v_{k}$ functions, selecting a specific $v\in\ker W$ as a common ground.
@@
A possible approach to overcome the just-discussed limitations of Eq.˜11 would be to use variational methods (like in Eq.˜2) to compute the sub-reconstructions. These non-linear methods are known to remove or strongly reduce these long-range artifacts. However, as discussed in [22], their non-linear nature contrasts with the assumptions of Eq.˜11. This may then result in a degraded reconstruction quality.
@@ 3 Results
We now analyze N2G’s performance against existing unsupervised methods for reconstructing both synthetic and real data. The reference methods are LS, TV, the GIDC algorithm from [17] (based on DIP), and INRs. For the synthetic test cases, we first derive a method to module the noise and compare different noise levels.
@@ 3.1 Synthetic data generation & quantification
We first evaluate the performance of the proposed method on synthetic data, as it provides the ground truth and precisely allows us to control the noise added to the data.
We extracted a $100\times 100$ pixels image to be our ground truth (phantom, e.g., from Figs.˜5f and 6f) from [23], a curated collection of chromosome images. We generated synthetic random masks with a normalized half-Gaussian distribution and forward projected the phantom to create the corresponding buckets. We added Poisson noise to the buckets by first multiplying them by the maximum expected emitted number of photons per pixel. We fed the resulting mean expected photons per realizations to the Poisson distribution function. We then normalized the signals by dividing the noisy buckets by the same number of expected photons, resulting in signals:
@@ 3.2 Synthetic data reconstructions across various noise levels
For the TV, GIDC, INR, and N2G reconstructions, the regularization weight $\lambda$ was chosen by cross-validation, by putting aside a set of realizations ($10\%$ of the total) and using them to assess the best value over a wide range [21]. For TV, we use a $\lambda\in[10^{-5},10^{-1}]$, while for GIDC, INR, and N2G we used a $\lambda\in[10^{-7},10^{-4}]$. This difference in range is due to the different data normalization used by the different algorithm implementations. As regularization $R$ for GIDC, INR and N2G, we also used the TV term. Although stronger and more advanced priors could have been used, we wanted a direct comparison against the well-known and used TV-min-based variational regularization. Moreover, the TV term has very simple and widely understood implementation and behavior. To improve the procedure’s robustness, we ran it three times with different cross-validation sets for each run and averaged the cross-validation loss values between those different runs. For the INR reconstructions, we used a 3-layer feed-forward NN with 512 neurons per layer, sinusoidal activation functions [18], and 256 Fourier embeddings [26]. Regarding specifically GIDC and N2G, we used a U-net [14] with 24 features and 3 decomposition levels ($\sim$ 276k parameters). We learned the U-net and INR weights with 3000 and 7000 epochs of the Adam optimizer [27] respectively, with a learning rate of $2\times 10^{-4}$ and weight decay of $1\times 10^{-2}$ for both architectures. We selected the model weights corresponding to the epoch in which the cross-validation loss function is minimized. For N2G, we used 4 partitions and 6 permutations ($K=4$, and $P=6$). These hyperparameter values were also selected cross-validation over a grid search.
@@
On the left side of Fig.˜3a, we present the peak signal-to-noise ratio (PSNR) of the reconstructions against the phantom. It is evident that N2G consistently achieves superior reconstructions, with PSNRs surpassing those of competing methods by 1–2 dB, highlighting its robustness across diverse imaging conditions. Nonetheless, its advantage seems to diminish under low-photon scenarios. While counterintuitive, this is easily explained by the fact that the lower frequencies are less affected by Poisson noise. PSNR takes all frequencies into account, resulting in a bias towards lower frequencies with low photon counts. The SSIM, by construction, is instead less affected by lower frequencies. For this reason, on the right side of Fig.˜3a, we see that N2G has a growing lead in SSIM for lower photon counts. We also observe that for high photon numbers (equating to low noise), the compression ratio becomes the limiting factor for reconstruction accuracy, and both PSNR and SSIM reach a plateau. From Fig.˜3b, we have instead a perspective on how the resolution is degraded with increasing noise levels. We observe that N2G better preserves the reconstruction resolution, by showing a drop-off at lower photon counts than other methods.

### v3

@@ Abstract
We present a new self-supervised deep-learning-based Ghost Imaging (GI) reconstruction method, which provides unparalleled reconstruction quality for noisy acquisitions among unsupervised methods. We present the supporting mathematical framework and results from theoretical and real data use cases. Self-supervision removes the need for clean reference data while offering strong noise reduction. This provides the necessary tools for addressing signal-to-noise ratio concerns for GI acquisitions in emerging and cutting-edge low-light GI scenarios. Notable examples include micro- and nano-scale x-ray emission imaging, e.g., x-ray fluorescence imaging of dose-sensitive samples. Their applications include in-vivo and in-operando case studies for biological samples and batteries.
journal: opticajournalarticletype: Research Article
@@ 1 Introduction
Thanks to all these advantages, GI has the potential to spark many breakthroughs in applications in various fields ranging from biomedical imaging to remote sensing. However, GI has only seen limited applicability so far, especially in the above-mentioned nano-scale imaging applications. In these cases, the observed signals are affected by overwhelmingly high Poisson noise levels due to their very small photon fluxes. This noise type does not enjoy the intrinsic noise reduction of GI, known as Felgett’s advantage, thus reducing GI’s advantage over PB [8]. Both traditional methods, based on regularized variational reconstructions (e.g., Total Variation minimization - TV-min [10]), and modern machine learning (ML) techniques (both supervised [11] and unsupervised [12]) are not specifically designed for this scenario. Here, we propose a GI reconstruction method particularly suited for working with acquisitions affected by high noise levels, regardless of whether the noise type is Gaussian or Poisson. Our method leverages the latest developments in unsupervised ML techniques, so it does not require high-quality reference data. This work is a step towards practical applications of GI, particularly for photon-limited scenarios like nano-scale x-ray emission imaging [13].
@@ 2.1 Existing methods
GI’s ability to obtain images with more pixels than the acquired realizations results in an undersampled reconstruction problem. These reconstructions require some prior knowledge (regularization) to recover accurate images: Unregularized least-squares (LS) reconstructions present specific high-frequency large-scale artifacts that resemble high noise levels. Methods employing convolutional neural networks (CNNs) are currently the best-performing tools for image processing applications. They can capture small- and large-scale correlations and learn the most suitable representations to capture the features of the treated signals (acting like regularizers) [15]. Thus, they are also better suited for GI reconstructions than traditional variational methods, e.g., TV-min, which use strong but also rather simple a priori assumptions on the types of features present in the reconstructed images.
Table 1: Comparison of GI Reconstruction Methods
Method Description Advantages Limitations
Unregularized Least-Squares (LS) Direct reconstruction without regularization Simple, no prior assumptions High-frequency artifacts; high noise levels; no compression
Regularized Variational (e.g., TV) Uses strong a priori assumptions (e.g., sparsity, smoothness) Robust; cheap; reference in literature Oversimplified assumptions; limited feature capture
Supervised denoising (e.g., DeepGhost) Uses pairs of LS and high-quality reconstructions Effective for known image types Dependent on large, similar training data; sensitive to parameter variations
Self-Supervised denoising (e.g., N2V, N2N) No pre-training; denoising in image space No clean data required Unable to deal with missing realization artifacts
Unsupervised (e.g., GIDC, INRs) No pre-training; physics-informed learning No clean data required; best results with high compression Reduced quality with high noise; relies on regularization/early stopping
@@
Both ML methods and CNN architectures have seen an explosion in development and applications over the last decade. Quite a few methods have been specifically developed for GI reconstructions. Notable examples are supervised methods [11], where a CNN (e.g., a U-net [16]) compares LS GI reconstructions from undersampled acquisitions against their corresponding high-quality reconstructions. The network learns to identify and correct the corrupted features in the LS images. This approach heavily depends on large amounts of known examples that present strong similarity to the images of interest. It also requires specific training data for each acquisition parameter variation (e.g., noise levels, image size, levels of undersampling, etc).
@@
Untrained generator-network methods do not require pre-training against high-quality reference images, thereby removing the need for large databases [17, 18]. They can incorporate knowledge of the image formation model in the learning algorithm (e.g., physics-informed ML), and train the model to produce an image that fits the acquired data. A good example is the method called GIDC [12], which is based on the deep image prior (DIP) [15]. As a drawback, these methods exhibit lower reconstruction quality for high noise levels in the acquired data. They delegate noise reduction solely to regularization and early stopping. Similar to DIP-based methods, implicit neural representations (INRs) -based methods use untrained multi-layer perceptron (MLP) NNs to reconstruct an image without high-quality reference data [19, 13]. They accept pixel coordinates instead of input images. During training, the MLP learns a representation of the imaged object, which produces the desired image for the input coordinates.
@@
Self-supervised methods like Noise2Noise (N2N) use the same training procedure as supervised methods, but they use noisy data as target [20]. They only require two noisy realizations of the same measurement (e.g., two noisy images of the same scene) and exploit the fact that the only difference between the two should be the noise. For each of the two measurements, they train the CNN to predict the other. As previously mentioned, CNNs can capture small and large-scale correlations, leading them to learn the features of the measurements while discarding their uncorrelated parts, i.e., the noise. Similarly, single-image self-supervised denoising methods have also emerged, including Noise2Void [21], and a few more recent developments [22, 23, 24, 25]. Compared to N2N, these methods only exploit the spatial correlation of imaging data. N2N can be adapted to work with inverse problems like computed tomography [26], while single-image methods can be directly applied to least-squares reconstructions. However, as we will see in Subsec.˜2.5, they do not handle the so-called missing realization problem, which refers to artifacts arising from not having enough realizations to represent the imaged objects, and that is usually an important source of noise in GI reconstructions. A brief comparison of all the mentioned methods is given in table˜1.
@@ 2.2 Forward model and variational methods
$\mathbb{E}_{y}\lVert WN_{\theta}(r)-y\rVert_{2}^{2}=\mathbb{E}_{b,\epsilon}\lVert WN_{\theta}(W^{\dagger}(b+\epsilon))-(b+\epsilon)\rVert_{2}^{2}$ (4)
@@ 2.3 Proposed method: Noise2Ghost (N2G)
We partition the realizations in $k$ splits $k\in[1,K]$, producing $K$ different sub-reconstructions (i.e., a k-tuple of reconstructions). Each sub-reconstruction in the said k-tuple has unique noise characteristics compared to the other reconstructions, and it provides different information on the FoV. From the partitioned data vectors $y_{k}$ we compute the sub-reconstructions $x_{k}$ as follows (Fig.˜2a and algorithm˜1):
@@
Algorithm 1 Split Realizations and Least-Squares Reconstructions
1:Realizations $W$, measurements $y$, number of splits $K$
2:Sub-reconstructions $\{x_{k}\}_{k=1}^{K}$
[…]
4:for $k=1$ to $K$ do
5:  $s_{k}\leftarrow s[M(k/K)+1,M((k+1)/K)]$ $\triangleright$ Slice $s$
6:  $W_{k},y_{k}\leftarrow W[s_{k}],y[s_{k}]$ $\triangleright$ Partition $W$ and $y$
7:  $x_{k}\leftarrow W_{k}^{\dagger}y_{k}$ $\triangleright$ Least-squares reconstruction
8:end for
9:return $\{x_{k}\}_{k=1}^{K}$, $\{W_{k},y_{k}\}_{k=1}^{K}$
Each sub-reconstruction $x_{k}$ represents the same signal but has a different noise. In N2G, we feed each of these sub-reconstructions to the model, and then we optimize the following minimization problem (Fig.˜2b and algorithm˜2) by comparing them against the realizations that were not used to reconstruct them:
@@
Since the different sub-reconstructions $x_{k}$ should represent the same signal, we would like the model to produce the same result for each. Thus, the model should learn to differentiate between the actual signal and random noise. This is in opposition to Eq.˜3a, where the same realizations were used for the input reconstruction and as targets in the learning loss. Finally, Fig.˜2c presents how Eq.˜6b produces the reconstructed images
Algorithm 2 N2G Training Loop
1:Sub-reconstructions $\{x_{k}\}_{k=1}^{K}$, realizations $\{W_{k},y_{k}\}_{k=1}^{K}$, regularization $R$, parameter $\lambda$
2:Trained model parameters $\hat{\theta}$
3:Initialize model $N_{\theta}$ randomly
4:for each epoch do
5:  for $k=1$ to $K$ do
6:   for $i=1$ to $K$, $i\neq k$ do
7:     $\mathcal{L}_{k,i}=\|W_{i}N_{\theta}(x_{k})-y_{i}\|_{2}^{2}$ $\triangleright$ Self-supervised losses
8:   end for
9:   $\mathcal{L}_{\text{TV},k}=R(N_{\theta}(x_{k}))$ $\triangleright$ Prior (TV) losses
10:   $\mathcal{L}_{k}\leftarrow(\frac{1}{2}\sum_{i\neq k}\mathcal{L}_{k,i})+\lambda\mathcal{L}_{\text{TV},k}$ $\triangleright$ Sub-reconstruction losses
11:  end for
12:  $\theta\leftarrow\theta-\nabla_{\theta}(\sum_{k}\mathcal{L}_{k})$ $\triangleright$ Update parameters
13:end for
14:return $\hat{\theta}$
@@ 2.4 Self-supervised random noise reduction
$\mathbb{E}_{b,\epsilon,i\neq k}$ $\lVert W_{i}N_{\theta}(x+v_{k}+W^{\dagger}\epsilon_{k})-b_{i}\rVert_{2}^{2}+\mathbb{E}_{b,\epsilon,i\neq k}\lVert b_{i}-(b_{i}-\epsilon_{i})\rVert_{2}^{2}$ (9a)
$=\mathbb{E}_{b,i\neq k}$ $\lVert W_{i}N_{\theta}(x_{k})-b_{i}\rVert_{2}^{2}+\mathbb{E}_{\epsilon,i\neq k}\lVert\epsilon_{i}\rVert_{2}^{2}$ (9b)
@@ 2.5 Relationship with Noise2Inverse
where $x_{i}=W_{i}^{\dagger}y_{i}$ is the pseudo-inverse (LS reconstruction) of the partial set of masks $W_{i}$ and corresponding buckets $y_{i}$ with $i\in[1,K]$. Theoretical grounds support this formulation, and it should retrieve the same result as Eq.˜6, regarding its noise reduction aspect. However, it does not address the missing realization artifacts and suffers from strong technical limitations. In ghost imaging, we aim at reducing the number of realizations as much as possible, which results in faster acquisition time and potentially reduced dose deposition. This exacerbates the rise of related artifacts. Missing realization and random noise artifacts are strongly structured and have long-range correlations, i.e., they generally cover the whole FoV. Commonly used models like CNNs have a limited receptive field (i.e., the region around each pixel that the CNN probes and that provides the context for the said pixel) that may not extend to the entire FoV. Depending on the model, this could have a strong negative impact on the denoising performance.
More precisely, about the missing realization artifacts, in [26] it is stated that Eq.˜10 is unable to cope with them. The LS sub-reconstructions cope with the missing realizations by acting as if their bucket values were set to 0. This selects one specific solution $v_{k}$ for each $x_{k}$, out of the infinite solutions in $\ker W_{k}$ that satisfy Eq.˜1. This specific choice of $v_{k}$ is probably incorrect, but is imposed by N2I as the learning target in Eq.˜10. The model learns to discard the inconsistent features of different $v_{k}$ functions, selecting a specific $v\in\ker W$ as a common ground.
@@
A possible approach to overcome the just-discussed limitations of Eq.˜10 would be to use variational methods (like in Eq.˜2) to compute the sub-reconstructions. These non-linear methods are known to remove or strongly reduce these long-range artifacts. However, as discussed in [26], their non-linear nature contrasts with the assumptions of Eq.˜10. This may then result in a degraded reconstruction quality.
@@ 2.6 Data augmentation
2.7 Models
Here, we briefly discuss the machine learning models, in particular neural networks (NN), that N2G could leverage. The model should accept multiple images of the same object as input. This is incompatible with INRs, which learn coordinate-based representations of the signal and thus take coordinates as inputs. CNNs and other MLP-based NNs are instead good candidates. Both NN types take images as input and can learn short- and long-range pixel correlations. Thus, N2G is not tied to a specific NN architecture and can take advantage of future developments in that area. MLP-based NNs examples include the Visual Transformer (ViT) [28] and the swin-transformer [29]. CNNs examples include U-Net [16], DnCNN [30], and MS-D net [31]. Regarding the model size, one should try to work with small NNs to minimize the computation time and resources and to limit overfitting. For instance, we suggest models with parameter numbers in the order of 100k for a $100\times 100$ pixels image. However, some models like MS-D net could achieve similar results with 10-20k parameters for similar image sizes. In our tests, we had the best results with U-Net and DnCNN with around 100-300k parameters. Further details will be provided with the results.
2.8 Computational Performance & Requirements
N2G presents specific computational and memory demands that require to be considered. It requires a $KP$-fold increase in operations, compared to GIDC (e.g., $\times 24$ for the tests performed in Sec.˜3). This does not correspond to a proportional increase in computational time, thanks to batching and GPU parallelism, as detailed in the Supplementary Information, Sec. 5. The actual increase in reconstruction time is approximately 2.1 to 2.5 times. On the other hand, memory usage increases proportionally, requiring $\times KP$ activation weights during backpropagation. This is critical for GPUs, where memory constraints are typically more stringent compared to CPU RAM. The increased memory requirements can limit the size of images and the extent of noise reduction that can be processed with N2G. To mitigate this, gradient accumulation can be used, where backpropagation weights are accumulated over smaller batches of model passes. While this approach can impact reconstruction performance, it effectively allows for the processing of larger reconstruction sizes. Looking ahead, this strategy could enable implementing multi-GPU reconstructions, further enhancing the capabilities of N2G. Furthermore, adjustments of “activation checkpointing” could further yield reductions in GPU memory consumption.
Note that, due to the intrinsic iterative nature of the methods discussed here, their use in real-time applications should be considered relative to the acquisition time. As presented in Sec. 5 of the Supplementary Information, reconstructions can take from several seconds to several minutes. This is incompatible with high-throughput acquisitions, where full datasets are produced short time-frames of less than a second or millisecond. However, given the sequential nature of GI acquisitions, this scenario is unlikely. In particular, for x-ray GI, the current acquisition time is on the order of a few hours per image [27]. In this context, N2G’s increased computational time requirements over the existing methods are negligible, even when performing a full hyperparameter grid search.
@@ 3 Results
We now analyze N2G’s performance against existing unsupervised methods for reconstructing both synthetic and real data. The reference methods are LS, TV, the GIDC algorithm from [12] (based on DIP), and INRs. For the synthetic test cases, we first derive a method to module the noise and compare different noise levels. Additionally, for the synthetic study cases, we also compare the results against N2V and N2I, two common self-supervised denoising methods.
@@ 3.1 Synthetic data generation & quantification
We first evaluate the performance of the proposed method on synthetic data, as it provides the ground truth and precisely allows us to control the noise added to the data. In Sec. 6 of the Supplementary Information, we present and analyze additional synthetic test cases, with increasing image complexity.
We extracted a $100\times 100$ pixels image to be our ground truth (phantom, e.g., from Figs.˜5h and 6h) from [32], a curated collection of chromosome images. We generated synthetic random masks with a normalized half-Gaussian distribution and forward projected the phantom to create the corresponding buckets. We added Poisson noise to the buckets by first multiplying them by the maximum expected emitted number of photons per pixel. We fed the resulting mean expected photons per realizations to the Poisson distribution function. We then normalized the signals by dividing the noisy buckets by the same number of expected photons, resulting in signals:
@@ 3.2 Synthetic data reconstructions across various noise levels
For the TV, GIDC, INR, and N2G reconstructions, the regularization weight $\lambda$ was chosen by cross-validation, by putting aside a set of realizations ($10\%$ of the total) and using them to assess the best value over a wide range [27]. For TV, we use a $\lambda\in[10^{-5},10^{-1}]$, while for GIDC, INR, and N2G we used a $\lambda\in[10^{-7},10^{-4}]$. This difference in range is due to the different data normalization used by the different algorithm implementations. As regularization $R$ for GIDC, INR and N2G, we also used the TV term. Although stronger and more advanced priors could have been used, we wanted a direct comparison against the well-known and used TV-min-based variational regularization. Moreover, the TV term has very simple and widely understood implementation and behavior. To improve the procedure’s robustness, we ran it three times with different cross-validation sets for each run and averaged the cross-validation loss values between those different runs. For the INR reconstructions, we used a 3-layer feed-forward NN with 512 neurons per layer, sinusoidal activation functions [19], and 256 Fourier embeddings [35]. Regarding specifically GIDC and N2G, we used a U-net [16] with 20 features and 3 decomposition levels ($\sim$ 192k parameters). We learned the U-net and INR weights with 5000 and 7000 epochs of the Adam optimizer [36] respectively, with a learning rate of $3\cdot 10^{-4}$ and weight decay of $1\cdot 10^{-2}$ for both architectures. We selected the model weights corresponding to the epoch in which the cross-validation loss function is minimized. For N2G, we used 4 partitions and 6 permutations ($K=4$, and $P=6$). For a more detailed discussion of the choice of the $K$ and $P$ parameters, we refer to Sec. 4 of the Supplementary Information.
@@
On the top left side of Fig.˜3, we present the peak signal-to-noise ratio (PSNR) of the reconstructions against the phantom. It is evident that N2G consistently achieves superior reconstructions, with PSNRs surpassing those of competing methods by 1–2 dB, highlighting its robustness across diverse imaging conditions. Nonetheless, its advantage seems to diminish under low-photon scenarios. While counterintuitive, this is easily explained by the fact that the lower frequencies are less affected by Poisson noise. PSNR takes all frequencies into account, resulting in a bias towards lower frequencies with low photon counts. The SSIM, by construction, is instead less affected by lower frequencies. For this reason, on the top right side of Fig.˜3, we see that N2G has a growing lead in SSIM for lower photon counts. We also observe that for high photon numbers (equating to low noise), the compression ratio becomes the limiting factor for reconstruction accuracy, and both PSNR and SSIM reach a plateau. From Fig.˜3, we have instead a perspective on how the resolution is degraded with increasing noise levels. We observe that N2G better preserves the reconstruction resolution, by showing a drop-off at lower photon counts than other methods.